What is the magnitude of the resultant of the forces F1=-3i+2j-k, F2=2i-j+3k and F3=3i+j-k measured in dynes ?


1
Expert's answer
2021-01-14T19:55:56-0500

"\vec{F_1}=-3i+2j-k"

"\vec{F_2}=2i-j+3k"

"\vec{F_1}=3i+j-k"

"\text{the resultant force is the vector sum of all forces}"

"\vec{F}=\vec{F_1}+\vec{F_2}+\vec{F_3}"

"\vec{F}= -3i+2j-k+2i-j+3k+3i+j-k"

"\vec{F}=2i+2j+k"

"|\vec{F}|=\sqrt{2^2+2^2+1}=3"

"\text{the default force is measured in newtons}"

"1J=10^5dyn"

Answer:"\vec{F}=2i+2j+k \ ;|\vec{F}|=3*10^5dyn"